OpenStudy (anonymous):

How would I factor this: x^3 + 3x^2 – 10x – 24 If it's possible please add all the steps. I already know the answer I just don't know how to get it

rvc (rvc):

division synthetic

OpenStudy (anonymous):

Thank you rvc

OpenStudy (anonymous):

x^3 + 3x^2 – 10x – 24 =x^3 +2x^2 +x^2+2x-12x-24 = x^2(x+2) +x(x+2) -12(x+2) =(x+2)(x^2+x-12) =(x+2)(x^2+4x-3x-12) =(x+2)(x(x+4)- 3(x+4)) =(x+2)(x+4)(x-3)

OpenStudy (anonymous):

Let f(x)=x^3+3x^2-10x-24 Therefore f(1)=1^3+3*1^2-10*1-24=-30 f(-1)=.... f(2)=...... f(-2)=0 now we'll stop here....because we have got 0 since f(-2)=0,therefore x+2 is a factor of f(x) nw we'll divide f(x) by x+2......we'll get x^2+x-12 Therefore f(x)=(x+2)(x^2+x-12)=(x+2)(x-3)(x+4)